# Principal component analysis balancing prediction and approximation accuracy for spatial data

### v2

@@ 6 Discussion
The procedure described in Algorithm 1 depends on a specified number $r$ of PCs. Since increasing $r$ will reduce the total MSRE and increase the total MSPE, one could examine the trajectory of MSPE and MSRE with respect to $r$, and identify the ‘elbow’ of these curves as a practical choice of $r$; this reflects the threshold beyond which adding more PCs would not alter either the prediction or representation meaningfully. Practitioners may also be interested in estimating the variability of evaluation metrics such as TMSE, MSPE and MSRE to better understand the trade-off between them. While our method does not directly provide uncertainty estimates, one could conduct bootstrap to assess this, with the potential caveat of sample splitting on correlated data.
@@ Appendix A Proof
$\frac{\partial f}{\partial\eta}=2\gamma Z^{\top}(Z\eta-{\mathcal{S}}\mathcal{D}q)+2\lambda_{1}P\eta.$
Since the Hessian $\partial^{2}f/\partial\eta^{2}=2\gamma Z^{\top}Z$ is positive semidefinite, we can profile out $\eta$ by setting it to the minimizer $\tilde{\eta}(q):=(\gamma Z^{\top}Z+\lambda_{1}P)^{-1}(\gamma Z^{\top}{\mathcal{S}}\mathcal{D}q)$. The objective function can thus be rearranged as

### v3

@@ Appendix A Proof
$\frac{\partial f}{\partial\eta}=2\gamma Z^{\top}(Z\eta-{\mathcal{S}}\mathcal{D}q)+2\lambda_{1}P^{1/2}\eta,$
where the square root $P^{1/2}$ is well-defined because $\tilde{K}$ and $\tilde{Q}$ are both positive semidefinite. Since the Hessian $\partial^{2}f/\partial\eta^{2}=2\gamma Z^{\top}Z+2\lambda_{1}P^{1/2}$ is positive semidefinite, we can profile out $\eta$ by setting it to the minimizer $\tilde{\eta}(q):=(\gamma Z^{\top}Z+\lambda_{1}P)^{-1}(\gamma Z^{\top}{\mathcal{S}}\mathcal{D}q)$. The objective function can thus be rearranged as
@@ B.3 Additional Data Analysis Results
Tables S1 and S2 report the standard deviations of all evaluation metrics in Tables 1 and 2 respectively, capturing their variability across 10 cross-validation folds.
The larger variability of MSRE-trn for RapPCA is expected. We recall that RapPCA includes classical PCA as a special case; and therefore for each PC, RapPCA may revert to classical PCA and have smaller MSRE-trn in some cross-validation folds than others. Nevertheless, the comparable variability of TMSE between RapPCA and other methods indicate that the overall performance and predictability-representability tradeoff of RapPCA remain consistently superior to other benchmark methods.
Overall Metrics Individual MSEs
TMSE MSPE MSRE-trn PC1 PC2 PC3
PCA 4.45 3.50 0.16 2.24 1.60 0.52
PredPCA 4.80 3.22 0.30 1.95 1.49 0.39
RapPCA 4.26 2.73 0.65 2.25 1.51 0.69
Table S1: Standard deviations for all evaluation metrics across 10 cross-validation folds for the Seattle traffic-related air pollution data.
Overall Metrics Individual MSEs
TMSE MSPE MSRE-trn PC1 PC2 PC3
PCA 10.86 3.66 0.90 2.22 1.79 0.76
PredPCA 11.75 2.96 1.39 1.06 1.13 0.74
SpatialPCA 11.45 4.74 1.21 2.24 2.13 0.90
SpatialPCA: built-in 12.04 5.22 1.21 2.18 2.06 0.96
RapPCA 10.82 4.11 6.04 2.21 1.59 1.16
Table S2: Standard deviations for all evaluation metrics across 10 cross-validation folds for the breast tumor data.